if the angle between the pass axis of polariser and analyser is 45 degree write the ratio of intensities  of orginal light and the transmitted light after passing through the analyser

Grade:12

2 Answers

Biswajit Das
9 years ago
Let the intensity of the original light is I0.

In acoordance to Malu’s law, the intensity of transmitted light after passing through the analyser will be,

I (θ) = I0cos2θ
Hereθ =45°
So,I (θ) = I0cos2(45°) = I0/2
Thus, I0/I = 2
From the above observation we conclude that,the ratio of intensities of orginal light and the transmitted light after passing through the analyser will be 2.
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Let the intensity of the original light is I0.
In acoordance to Malu’s law, the intensity of transmitted light after passing through the analyser will be,
I (θ) = I0cos2θ
Hereθ = 45°
So,I (θ) = I0cos2(45°)
= I0/2
Thus, I0/I = 2
From
the above observation we conclude that,the ratio of intensities of
orginal light and the transmitted light after passing through the
analyser will be 2.

Thanks and Regards